OpenStudy (anonymous):

Find all solutions of the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to two decimal places where appropriate.) 7 cos θ + 1 = 0

OpenStudy (tkhunny):

Have you considered solving for \(\cos(\theta)\)?

OpenStudy (anonymous):

Yeah I got .14+2kpi, but I need another answer. I'm not sure how to find the second half.

OpenStudy (tylerd):



OpenStudy (tkhunny):

\(\cos(\theta) = \cos(2\pi - \theta)\). It's like a thing.

OpenStudy (tylerd):

pi-1.4 radians pi+1.4radians

OpenStudy (tylerd):

the adjacent side is negative. so your dealing with quadrants 2 and 3